OpenStudy (anonymous):

Can someone explain to me how to solve this? I dont understand how to do it? I have to verify the identity. sec^2 x = 1 + tan^2 x

OpenStudy (irishboy123):

Pythagoreas' theorem draw the triangle and match it all up !!!!

OpenStudy (irishboy123):

....with the definitions of sec and tan, of course

OpenStudy (anonymous):

@Directrix so then would it break up into the (sin^2x)/(sin^2x)+(cos^2x)/(sin^2x)=1/(sin^2x) equation?

OpenStudy (anonymous):

@Directrix i did study it but i just dont completely understand how to solve it or work it out

OpenStudy (anonymous):

i have studied the sin^2(theta)+cos^2(theta)=1 but i dont understand how to verify the identity

OpenStudy (irishboy123):

Pythagoreas' theorem

OpenStudy (anonymous):

@IrishBoy123 so then to verify it do i make it (sin^2x)/(sin^2x)+(cos^2x)/(sin^2x)=1/(sin^2x) ?

OpenStudy (irishboy123):

try it the other way round you have mjust proved that \(1 + \cot^2 x = \sec^2 x\)

OpenStudy (irishboy123):

which means "you're doing good"

OpenStudy (anonymous):

what do you mean by the other way around? @IrishBoy123

Directrix (directrix):

sin^2 x + cos^2 x = 1 Divide each term by cos^2 x sin^2 x + cos^2 x = 1 ----- ------- ----- cos^2 x cos^2 x cos^2 x tan^2 x + 1 = sec^2 x You need to know the basic trig identities for this to make sense. @Yessy16c